# Certifiably Optimal Anisotropic Rotation Averaging

### v1

@@ C.1 Synthetic experiments
Figure C.2: Histograms of rotation errors wrt. ground truth (rows 1,3,5). Corresponding solver runtime (s) wrt. the increasing number of cameras (rows 2,4,6). Results are shown for different fractions $p$ of observed relative rotations.
We also present complementing results of the synthetic experiments of Sec 5.1 in the paper. Figure C.2 shows the rotation error histograms and the dependency of the runtime on the number of cameras for the other fractions of the observed relative rotations. These results are in line with the analysis provided in the paper — using the proposed objective leads to lower errors and using the proposed constraints speeds up the employed SDP solver.
C.2 Real experiments
@@
$\sqrt{\sum_{i}\min\{\Delta\bm{\omega}_{i}^{-\top}H_{i}\Delta\bm{\omega}_{i}^{-},\Delta\bm{\omega}_{i}^{+\top}H_{i}\Delta\bm{\omega}_{i}^{+}\}},$ (C.13)
where the minimization is done to account for the sign ambiguity in the axis-angle representation. We also present the RMS angular errors for a better geometric interpretation. As shown in Table C.1, in many cases, the proposed method leads to a much lower error.
@@
LU Sphinx SDP-O(3)-iso 0.3879 0.46
SDP-cSO(3) 0.2067 0.36
Round Church SDP-O(3)-iso 0.6308 0.59
SDP-cSO(3) 0.3683 0.54
UWO SDP-O(3)-iso 1.4814 1.19
SDP-cSO(3) 0.7267 0.86
Tsar Nikolai I SDP-O(3)-iso 0.6871 0.48
SDP-cSO(3) 0.1878 0.22
Vercingetorix SDP-O(3)-iso 0.4312 1.53
SDP-cSO(3) 0.4228 1.42
Eglise Du Dome SDP-O(3)-iso 0.2239 0.24
SDP-cSO(3) 0.1882 0.21
King’s College SDP-O(3)-iso 0.2288 0.76
SDP-cSO(3) 0.1303 0.37
Kronan SDP-O(3)-iso 0.7380 0.76
SDP-cSO(3) 1.1114 1.38
Alcatraz SDP-O(3)-iso 1.3329 0.62
SDP-cSO(3) 1.0110 0.45
Museum Barcelona SDP-O(3)-iso 2.7101 0.79
SDP-cSO(3) 1.2164 0.46
Temple Singapore SDP-O(3)-iso 2.4197 0.86
SDP-cSO(3) 1.0760 0.55
Table C.1: Mahalanobis distance between the axis-angle vectors of the estimated and ground truth rotations and RMS angular errors (degrees) evaluated on the real datasets.

### v2

@@ D.1 Synthetic experiments
Figure D.2: Histograms of rotation errors wrt. ground truth (rows 1,3,5). Corresponding solver runtime (s) wrt. the increasing number of cameras (rows 2,4,6). Results are shown for different fractions $p$ of observed relative rotations.
We also present complementing results of the synthetic experiments of Sec 5.1 in the paper. Figure D.2 shows the rotation error histograms and the dependency of the runtime on the number of cameras for the other fractions of the observed relative rotations. These results are in line with the analysis provided in the paper — using the proposed objective leads to lower errors and using the proposed constraints speeds up the employed SDP solver.
D.2 Real experiments
@@
$\sqrt{\sum_{i}\min\{\Delta\bm{\omega}_{i}^{-\top}H_{i}\Delta\bm{\omega}_{i}^{-},\Delta\bm{\omega}_{i}^{+\top}H_{i}\Delta\bm{\omega}_{i}^{+}\}},$ (D.17)
where the minimization is done to account for the sign ambiguity in the axis-angle representation. We also present the RMS angular errors for a better geometric interpretation. As shown in Table D.1, in many cases, the proposed method leads to a much lower error.
@@
LU Sphinx SDP-O(3)-iso 0.388 0.46
Spectral 0.420 1.17
SDP-cSO(3) 0.207 0.36
Round Church SDP-O(3)-iso 0.631 0.59
Spectral 0.437 1.20
SDP-cSO(3) 0.368 0.54
UWO SDP-O(3)-iso 1.481 1.19
Spectral 6.125 7.07
SDP-cSO(3) 0.727 0.86
Tsar Nikolai I SDP-O(3)-iso 0.687 0.48
Spectral 0.344 0.71
SDP-cSO(3) 0.188 0.22
Vercingetorix SDP-O(3)-iso 0.431 1.53
Spectral 30.970 86.94
SDP-cSO(3) 0.423 1.42
Eglise Du Dome SDP-O(3)-iso 0.224 0.24
Spectral 0.119 0.22
SDP-cSO(3) 0.188 0.21
King’s College SDP-O(3)-iso 0.229 0.76
Spectral 0.251 1.00
SDP-cSO(3) 0.130 0.37
Kronan SDP-O(3)-iso 0.738 0.76
Spectral 2.622 4.36
SDP-cSO(3) 1.111 1.38
Alcatraz SDP-O(3)-iso 1.333 0.62
Spectral 0.667 0.80
SDP-cSO(3) 1.011 0.45
Museum Barcelona SDP-O(3)-iso 2.710 0.79
Spectral 16.588 7.35
SDP-cSO(3) 1.216 0.46
Temple Singapore SDP-O(3)-iso 2.420 0.86
Spectral 0.719 0.46
SDP-cSO(3) 1.076 0.55
Table D.1: Mahalanobis distance between the axis-angle vectors of the estimated and ground truth rotations and RMS angular errors (degrees) evaluated on the real datasets.